# IFT-UAM/CSIC-24-89 Insights into Dark Matter Direct Detection Experiments: Decision Trees versus Deep Learning

### v1

@@
cInstituto de Física Teórica, IFT-UAM/CSIC, 28049 Madrid, Spain
dDepartamento de Física Teórica, Universidad Autónoma de Madrid, 28049 Madrid, Spain
@@
The detection of Dark Matter (DM) remains a significant challenge in particle physics. This study exploits advanced machine learning models to improve detection capabilities of liquid xenon time projection chamber experiments, utilizing state-of-the-art transformers alongside traditional methods like Multilayer Perceptrons and Convolutional Neural Networks. We evaluate various data representations and find that simplified feature representations, particularly corrected S1 and S2 signals, retain critical information for classification. Our results show that while transformers offer promising performance, simpler models like XGBoost can achieve comparable results with optimal data representations. We also derive exclusion limits in the cross-section versus DM mass parameter space, showing minimal differences between XGBoost and the best performing deep learning models. The comparative analysis of different machine learning approaches provides a valuable reference for future experiments by guiding the choice of models and data representations to maximize detection capabilities.
@@ 1 Introduction
To distinguish a WIMP signal from the background is crucial, and new machine learning (ML) techniques could play a essential role in the near future. In fact, ML is showing promising results (see e.g. [27, 28, 29, 30, 31, 32, 33, 34]) in the field of particle physics and astroparticle physics. In particular, in DM direct detection phenomenology both supervised and semi-supervised approaches for direct detection were discussed in [35, 36, 37]. Several collaborations, like LUX [38] and CRESST [39], also have applied machine learning techniques to their data analysis. In [40, 41] was suggested to improve ANAIS-112 sensitivity to DAMA/LIBRA signal using machine learning techniques. In the case of directional dark matter experiments, DRIFT collaboration release [42], where the improvement of experiment sensitivity by the use of machine learning techniques was discussed. Additionally, the application of deep learning to distinguish between DM signals vs several backgrounds in the context of the proposed NEWSdm directional experiment was discussed in [43].
@@
This paper is organized as follows. In Section 2 we briefly describe the used experiments, proposed WIMP events and the types of backgrounds. Section 3 is devoted to the process of simulating events. We also discuss the different data representations that we compare. The architecture used, the training and testing procedure, are explained in Section 4. In Section 5, we discuss the results. Finally, in Section 6 we give the conclusion of this work.
@@ 3.1 Data Representation
Raw data: images
Within the first group, we consider the response of the detectors as images, as analysed in [35, 36]. In Fig. 1 we show two examples, the left image corresponds to a signal event with GeV, cm2, and the right image to an ER background event. Notice that both images are very similar (we have removed the axis labels to ease image classification, but signal and background events have also similar magnitudes). Each event is characterized by an image composed of four panels, the first two correspond to the responses or hit patterns of the bottom (upper left panel) and top (upper right panel) PMTs, where the relative position of the PMTs is shown by a circle and the intensity of the signal detected by each PMT is represented by a colour gradient. Additionally, the lower panels correspond to the largest S1 (left) and S2 (right) peaks as a function of time. The main advantage of this representation is to make explicit the correlation between the responses of each PMT that could be exploited by a ML algorithm like a CNN. On the other hand, one of the main drawbacks of working with images is that we need to generate a large dataset, which can be computationally very expensive to analyze and to store. We compressed the images generated with the simulator and tested different resolutions , , and pixels, and found no significant differences between them, therefore we present the results using pixels images.
Raw data: arrays
The second type of raw data that we consider also includes the same four detector responses, but instead of working with images, we store four vectors. The first one corresponds to the intensity of the bottom hit pattern, i.e. a number in the interval representing the intensity of each of the 121 PMTs, the second one with the top hit pattern, a vector with the intensity of the 127 PMTs, the third vector with the intensity of the largest S1 peak normalized to 1 as a function of time, and the fourth to the largest S2 peak. In this representation the third and fourth vectors have different length for every event because the program takes a measurement of the intensity (sample time) every 10 ns while the time duration of each signal can vary. To be able to analyse this information with ML we add extra zero values, a method called zero padding, to obtain vectors of the same length (we apply this procedure to the S1 and S2 peaks separately).
Processed data: S1 and S2 peak reduced features
Regarding the processed data, we construct a set of reduced features that are relevant shape-related quantities describing the S1 and S2 peaks, such as integrated areas and lengths. Some of the features considered in this work are based on the ones taken in Ref. [37]. The set of shape-related quantities has the following 14 quantities
@@
We also include an additional feature,
@@
Notice that the information about the time separation between pulses is not available in either of the previous raw data representation by construction. Therefore, we will combine with the raw data to be able to compare datasets fairly.
In Fig. 2 we show the correlation of some of the features. As we will show in Section 5 these correspond to the most important features determined by XGBoost (the distributions are independent of the classifier) for the discrimination between background and signal with GeV. Since we are only interested in showing the distributions and correlations between variables, the same number of background and signal events are presented.
Processed data: cS1 and cS2
Finally, we also consider cS1 and cS2. As explained in Section 2, these high-level features are similar to the total integrated area of S1 and S2 pulses (also to the sum of the intensity of all the PMTs) but corrected by factors that take into account the geometric effects of the TPC, position-dependent inhomogeneous effects, light collection efficiency, the drift electron lifetime under the drift field. For more details about the modelling of these corrections, see Ref. [44]. In this work, we use the default values included in Ref. [45, 46, 47]. The information in the cS1-cS2 plane is the one used by the collaboration using liquid xenon experiments to set limits to the DM parameter space, and naturally require a huge effort to characterize the detector response and to model the correction factors. The main advantage of this representation is its simplicity that allows a relative fast analysis even with complex models and ease the interpretation of the results. One of the goals of this work is to compare this simple representation under the scrutiny of state-of-the-art ML algorithms. Comparing the discrimination power of the resumed features and the raw data representations, we can check if the cS1-cS2 plane retains all the useful information. Moreover, this comparison also can be applied by the experimental collaborations to cross check the modelling of the S1 and S2 correction factors. In Fig. 3 we show an example with 1k signal and 1k background events in the cS1-cS2 plane, considering GeV.
@@ 4 Architecture and Training
For the transformer models, we divide the data into three subsets: 80% of the data is used for training, 10% for validation and the remaining 10% for testing. We use the Adam optimizer to train the models. The initial learning rate is set to , and to further refine the learning rate during training, we use a learning rate planner based on the plateau of the validation loss. If the validation loss does not improve for a certain number of epochs, the learning rate is reduced by a factor of two to allow for finer adjustments and to avoid possible local minima. Finally, when the learning rate reaches a value of , the training is aborted. In addition, we also performed a grid search to find the optimal set of hyperparameters that optimize the AUC.
@@
For further details about the algorithms employed, the code is available at [59].
@@ 5.1 Classifiers performance
| XGBoost | DNN | CNN | Transformer | ||
| Raw data | only raw | 0.919 | 0.924 | 0.918 | 0.93∗ |
| (images) | + | 0.942 | 0.958 | 0.951 | 0.965∗ |
| Raw data | only raw | 0.954 | 0.960 | 0.953 | 0.963† |
| (arrays) | + | 0.969 | 0.978 | 0.968 | 0.98† |
| Resumed | 14 features | 0.956 | 0.956 | 0.955 | 0.966‡ |
| features | + | 0.980 | 0.980 | 0.979 | 0.982‡ |
| 4 main + | 0.980 | 0.979 | 0.972 | 0.982‡ | |
| S1 vs S2 | cS1, cS2 | 0.980 | 0.980 | 0.980 | 0.980‡ |
In Table 2 we present one of the main results of our manuscript comparing the AUC, area under the receiver operating characteristic (ROC) curve, of different ML classifiers and data representations. The AUC is a common quantity to evaluate the performance of a binary classifier, where AUC=1 indicates a perfect classification and AUC=0.5 a completely random result.
@@
We can see in Table 2 that the worst result corresponds to the raw data representation with images (first row) with AUC. These results correspond to images with to ease the computation, but with , and pixels we found no significant differences. The main advantage of the images is to keep the spatial correlations, specially in the hit patterns of the top and bottom PMTs (see in Fig. 1 the non-trivial distribution of the PMTs). However, the possible benefit is out-weighted by the increased complexity of analysing an image, which implies an exponentially larger number of parameters to be tuned by the ML classifiers. Considering the two algorithms that are designed to take advantage of such spatial correlations, the CNN and transformer, the first one gives the lowest performance of all the method reported, while the more complex transformer is able to barely surpass the rest.
@@
Comparing the results using both raw data representations, images versus arrays, we notice an overall increase of the classifier performance when using the array representation. Recall that in the latter, one event is described with 4 arrays, two containing the response of each PMT (bottom and top) and two with the time series of the main pulses, i.e. the same information as in the image representation, but without the spatial correlation of the PMT. As explained before, this simpler representation not only allows for a significantly faster training with fewer parameters to tune but is more adequate, showing that the PMT spatial correlations do not provide significant information for the classification task. As before, the state-of-the-art transformer is the best performing algorithm, followed by the fully connected DNNs.
@@
We want to highlight that we have also studied different combinations between the raw data representations. Instead of considering a single image composed of four panels, like the ones shown in Fig. 1, we have analysed each panel independently, and also used them to train 4 different ML classifiers that we combine afterwards into a single fully connected layer to provide a single classification. Moreover, we replaced some panels with the array representation. For example, we used the images of the PMT responses to keep the spatial correlations combined with the arrays of the time series. To ease the interpretation of our results we do not show these possibilities, since there are lots of combinations and algorithm configurations, but no significant gain was achieved.
@@
Regarding the processed data representations, we can see that the 14 shape-related resumed features that describe the S1 and S2 pulses present a similar or slightly higher performance than the raw data as arrays, indicating that these set of variables capture the signal characteristics properly. Naturally, the smaller set of features implies smaller ML algorithms and faster training times.
@@
One feature that we have not included in the shape-related quantities is the time difference between the S1 and S2 signals, . If we add this feature, we notice a significant increase in performance. The same occurs for the raw data representations, where the information about the time difference between the pulses was not available. Looking only at the bottom right panel of Fig. 2, one might assume that provides no discrimination power and should not be included in the training procedure, as both its background and signal distributions are similar. However, the impact of becomes evident when examining its correlations with other features, especially those related to S2 (the signal originating from the ionization electrons drifting towards the upper part of the detector). For instance, in the second panel of the last row of Fig. 2, for pA S2 PE, events with s are more likely to be background. In contrast, when looking solely at the distributions of the pA S2 feature (second panel of the second row), we observe similar signal and background probabilities for pA S2 PE. Additionally, we have verified that the performance of a classifier trained with (pA S2, ) is significantly better than one trained with only (pA S2). This improvement is not observed for S1-related features; for example, the performance remains fairly similar whether using only (pA S1) or (pA S1, ).
@@
The relevance of can also be seen in Fig. 4 where we show the feature importance considering the gain metric for the XGBoost classifier. The most important features are both pA, the intensity of the signals, and in a close third place. The rest of the features seem not to be as relevant, however we noticed a significant performance drop if we only employ the 3 main features. We also need to include at least both pRMSW to keep a similar performance.
@@
Finally, in the last row of Table 2 we show the results for the classifiers using only two features: cS1 and cS2, two variables usually employed by the collaborations to perform their analyses. Recall that these quantities are related to shape-related resumed features pA, the intensity of the S1 and S2 signals but corrected by geometric and response factors taking into account the particular characteristics of the experiment. From our results we can see that the cS1-cS2 plane not only provides a simpler way to represent the data and ease the interpretation of the results, but also retains all the relevant information for the classification providing the highest AUC values. For that reasons cS1-cS2 is the best data representation. The only drawback is the determination of the correction factors, which implies a huge experimental effort. In that sense, the resumed feature representation can be useful to cross-check the modelling of those factors 111We have checked that AUC values obtain with cS1 and cS2 can not be obtained with only 3 resumed features: pA S1, pA S2 and . We need to including the 5 main features (according to the XGBoost gain metric) as can be seen comparing the last two rows of Table 2..
@@
Another important conclusion is that we have not found a significant difference in performance between the ML algorithms. Even more, the simple and very fast XGBoost method usually outperforms more complex and harder to tune classifiers like CNNs. Hence, in the following, we will just use XGBoost results. In Fig. 5 we show the ROC curves for some data representations and XGBoost, they correspond to ones with the lowest and highest AUC. The rest of the algorithms have similar ROC shapes, according to their AUC.

### v2

@@
cInstituto de Física Teórica, IFT-UAM/CSIC, Campus de Cantoblanco, 28049 Madrid, Spain
dDepartamento de Física Teórica, Universidad Autónoma de Madrid, Campus de Cantoblanco, 28049 Madrid, Spain
@@
The detection of Dark Matter (DM) remains a significant challenge in particle physics. This study exploits advanced machine learning models to improve detection capabilities of liquid xenon time projection chamber experiments, utilizing state-of-the-art transformers alongside traditional methods like Multilayer Perceptrons and Convolutional Neural Networks. We evaluate various data representations and find that simplified feature representations, particularly corrected S1 and S2 signals as well as a few shape-related features including the time difference between signals, retain critical information for classification. Our results show that while transformers offer promising performance, simpler models like XGBoost can achieve comparable results with optimal data representations. We also derive exclusion limits in the cross-section versus DM mass parameter space, showing minimal differences between XGBoost and the best performing deep learning models. The comparative analysis of different machine learning approaches provides a valuable reference for future experiments by guiding the choice of models and data representations to maximize detection capabilities.
@@ 1 Introduction
To distinguish a WIMP signal from the background is crucial, and new machine learning (ML) techniques could play a essential role in the near future. In fact, ML is showing promising results (see e.g. [27, 28, 29, 30, 31, 32, 33, 34]) in the field of particle physics and astroparticle physics. In particular, in DM direct detection phenomenology both supervised and semi-supervised approaches for direct detection were discussed in [35, 36, 37, 38]. Several collaborations, like LUX [39] and CRESST [40], also have applied machine learning techniques to their data analysis. In [41, 42] was suggested to improve ANAIS-112 sensitivity to DAMA/LIBRA signal using machine learning techniques. In the case of directional dark matter experiments, DRIFT collaboration release [43], where the improvement of experiment sensitivity by the use of machine learning techniques was discussed. Additionally, the application of deep learning to distinguish between DM signals vs several backgrounds in the context of the proposed NEWSdm directional experiment was discussed in [44].
@@
This paper is organized as follows. In Section 2 we briefly describe the used experiments, proposed WIMP events and the types of backgrounds. Section 3 is devoted to the process of simulating events. We also discuss the different data representations that we compare. The architectures used, the training and testing procedure, are explained in Section 4. In Section 5, we discuss the results. Finally, in Section 6 we give the conclusion of this work.
@@ 3.1 Data Representation
Raw data: images (dataset A)
Within the first group, we consider the response of the detectors as images, as analysed in [35, 36]. We call this representation dataset A1. In Fig. 1 we show two examples, the left image corresponds to a signal event with GeV, cm2, and the right image to an ER background event. Notice that both images are very similar (we have removed the axis labels to ease image classification, but signal and background events have also similar magnitudes). Each event is characterized by an image composed of four panels, the first two correspond to the responses or hit patterns of the bottom (upper left panel) and top (upper right panel) PMTs, where the relative position of the PMTs is shown by a circle and the intensity of the signal detected by each PMT is represented by a colour gradient. Additionally, the lower panels correspond to the largest S1 (left) and S2 (right) intensity peaks as a function of time, where the intensity of the signal corresponds to the total amount of light collected by all the PMTs. The main advantage of this representation is to make explicit the correlation between the responses of each PMT that could be exploited by a ML algorithm like a CNN. On the other hand, one of the main drawbacks of working with images is that we need to generate a large dataset, which can be computationally very expensive to analyze and to store. We compressed the images generated with the simulator and tested different resolutions , , and pixels, and found no significant differences between them, therefore we present the results using pixels images.
Raw data: arrays (dataset B)
The second type of raw data that we consider also includes the same four detector responses, but instead of working with images, we store four vectors per event and call this representation dataset B1. The first vector corresponds to the intensity of the bottom hit pattern, i.e. a number in the interval representing the intensity of each of the 121 PMTs, the second one with the top hit pattern, a vector with the intensity of the 127 PMTs, the third vector with the intensity of the largest S1 peak normalized to 1 as a function of time, and the fourth to the largest S2 peak. In this representation the third and fourth vectors have different length for every event because the program takes a measurement of the intensity (sample time) every 10 ns while the time duration of each signal can vary. To be able to analyse this information with ML we add extra zero values, a method called zero padding, to obtain vectors of the same length (we apply this procedure to the S1 and S2 peaks separately).
Processed data: S1 and S2 peak reduced features (dataset C)
Regarding the processed data, we construct a set of reduced features that are relevant shape-related quantities describing the S1 and S2 peaks, such as integrated areas and lengths. Some of the features considered in this work are based on the ones taken in Ref. [37]. The set of shape-related quantities, denoted dataset C1, has the following 14 quantities
@@
We also construct another set, called dataset C2, by including an additional feature,
@@
Notice that the information about the time separation between pulses is not available in either of the previous raw data representation by construction. Therefore, we will combine with both raw data representations, images and arrays, to be able to compare datasets fairly, and denote the respective datasets A2 and B2.
Furthermore, we consider a subset of the resumed features, called dataset C3, that corresponds to the five most important features determined by the XGBoost classifier for the discrimination between background and signal with GeV. As we will show in Section 5, this subset retains enough information to provide an excellent performance. In Fig. 2 we show the distributions and correlations (which are independent of the classifier) of the features in dataset C3, using the same number of background and signal events.
Processed data: cS1 and cS2 (dataset D)
Finally, we also consider cS1 and cS2 and denote this representation as dataset D. As explained in Section 2, these high-level features are similar to the total integrated area of S1 and S2 pulses (also to the sum of the intensity of all the PMTs) but corrected by factors that take into account the geometric effects of the TPC, position-dependent inhomogeneous effects, light collection efficiency, the drift electron lifetime under the drift field. For more details about the modelling of these corrections, see Ref. [45]. In this work, we use the default values included in Ref. [46, 47, 48]. The information in the cS1-cS2 plane is the one used by the collaboration using liquid xenon experiments to set limits to the DM parameter space, and naturally require a huge effort to characterize the detector response and to model the correction factors. The main advantage of this representation is its simplicity that allows a relative fast analysis even with complex models and ease the interpretation of the results. One of the goals of this work is to compare this simple representation under the scrutiny of state-of-the-art ML algorithms. Comparing the discrimination power of the resumed features and the raw data representations, we can check if the cS1-cS2 plane retains all the useful information. Moreover, this comparison also can be applied by the experimental collaborations to cross check the modelling of the S1 and S2 correction factors. In Fig. 3 we show an example with 1k signal and 1k background events in the cS1-cS2 plane, considering GeV.
@@ 4 Architecture and Training
For the transformer models, we divide the data into three subsets: 60 of the data is used for training, 15 for validation and the remaining 25 for testing. We use the Adam optimizer to train the models. The initial learning rate is set to , and to further refine the learning rate during training, we use a learning rate planner based on the plateau of the validation loss. If the validation loss does not improve for a certain number of epochs, the learning rate is reduced by a factor of two to allow for finer adjustments and to avoid possible local minima. Finally, when the learning rate reaches a value of , the training is aborted. In addition, we also performed a grid search to find the optimal set of hyperparameters that optimize the AUC.
@@ 5.1 Classifiers performance
XGBoost | DNN | CNN | Transformer | |||
A1 | Raw data | only raw | 0.919 | 0.924 | 0.918 | 0.930∗ |
A2 | (images) | + | 0.942 | 0.958 | 0.951 | 0.965∗ |
B1 | Raw data | only raw | 0.954 | 0.960 | 0.953 | 0.963† |
B2 | (arrays) | + | 0.969 | 0.978 | 0.968 | 0.980† |
C1 | Resumed | 14 features | 0.956 | 0.956 | 0.955 | 0.966‡ |
C2 | features | + | 0.980 | 0.980 | 0.979 | 0.982‡ |
C3 | 4 main + | 0.980 | 0.979 | 0.972 | 0.982‡ | |
D | cS1, cS2 | 0.980 | 0.980 | 0.980 | 0.980‡ | |
Uncertainty | 0.001 | 0.002 | 0.003 | 0.002 |
In Table 2 we present one of the main results of our manuscript comparing the AUC, area under the receiver operating characteristic (ROC) curve, of different ML classifiers and data representations. The AUC is a common quantity to evaluate the performance of a binary classifier, where AUC=1 indicates a perfect classification and AUC=0.5 a completely random result. To determine the AUC value and uncertainty, we have trained and tested each classifier 5 times using resampling with replacements, and report the mean and one standard deviation as final values. We found that the uncertainty is mainly associated with the ML method, therefore in Table 2 we show one error value per column.
@@
We can see in Table 2 that the worst results correspond to the dataset A1 (first row) with AUC. These results correspond to images with to ease the computation, but with , and pixels we found no significant differences. The main advantage of the images is to keep the spatial correlations, specially in the hit patterns of the top and bottom PMTs (see in Fig. 1 the non-trivial distribution of the PMTs). However, the possible benefit is out-weighted by the increased complexity of analysing an image, which implies an exponentially larger number of parameters to be tuned by the ML classifiers. Considering the two algorithms that are designed to take advantage of such spatial correlations, the CNN and transformer, the first one gives the lowest performance of all the method reported, while the more complex transformer is able to barely surpass the rest.
@@
Comparing the results using both raw data representations, dataset A1 (images) versus dataset B1 (arrays), we notice an overall increase of the classifier performance when using the array representation. Recall that in the latter, one event is described with 4 arrays, two containing the response of each PMT (bottom and top) and two with the time series of the main pulses, i.e. the same information as in the image representation, but without the spatial correlation of the PMT. As explained before, this simpler representation not only allows for a significantly faster training with fewer parameters to tune but is more adequate, showing that the PMT spatial correlations do not provide significant information for the classification task. As before, the state-of-the-art transformer is the best performing algorithm, followed by the fully connected DNNs.
@@
We want to highlight that we have also studied different combinations between the raw data representations. Instead of considering a single image composed of four panels, like the dataset A1 examples shown in Fig. 1, we have analysed each panel independently, and also used them to train 4 different ML classifiers that we combine afterwards into a single fully connected layer to provide a single classification. Moreover, we replaced some panels with the array representation, mixing datasets A1 and B1. For example, we used the images of the PMT responses to keep the spatial correlations combined with the arrays of the time series. To ease the interpretation of our results we do not show these possibilities, since there are lots of combinations and algorithm configurations, but no significant gain was achieved.
@@
Regarding the processed data representations, we can see that the 14 shape-related resumed features (dataset C1) that describe the S1 and S2 pulses present a similar or slightly higher performance than the raw data as arrays (dataset B1), indicating that these set of variables capture the signal characteristics properly. Naturally, the smaller set of features implies smaller ML algorithms and faster training times.
One feature that we have not discussed so far is the inclusion of the time difference between the S1 and S2 signals, . We notice a significant increase in performance when comparing datasets without and with this feature: A1 versus A2, B1 versus B2, and C1 versus C2, respectively.
@@
If we focus on the shape-related quantities, looking only at the bottom right panel of Fig. 2, one might assume that the addition of provides no extra discrimination power and should not be included in the training procedure, as both its background and signal distributions are similar. However, the impact of becomes evident when examining its correlations with other features, especially those related to S2 (the signal originating from the ionization electrons drifting towards the upper part of the detector). For instance, in the second panel of the last row of Fig. 2, for pA S2 PE, events with s are more likely to be background. In contrast, when looking solely at the distributions of the pA S2 feature (second panel of the second row), we observe similar signal and background probabilities for pA S2 PE. Additionally, we have verified that the performance of a classifier trained with (pA S2, ) is significantly better than one trained with only (pA S2). This improvement is not observed for S1-related features; for example, the performance remains fairly similar whether using only (pA S1) or (pA S1, ).
@@
The relevance of can also be seen in Fig. 4 where we show the feature importance considering the gain metric for the XGBoost classifier. The most important features are both pA, the intensity of the signals, and in a close third place. The rest of the features seem not to be as relevant, however we noticed a significant performance drop if we train a classifier only using the 3 main features. We also need to include at least both pRMSW (dataset C3) to keep a similar performance.
@@
Finally, in the second to last row of Table 2 we show the results for the classifiers using dataset D composed of only two features: cS1 and cS2, two variables usually employed by the collaborations to perform their analyses. Recall that these quantities are related to the shape-related resumed features pA, the intensity of the S1 and S2 signals but corrected by geometric and response factors taking into account the particular characteristics of the experiment. From our results we can see that dataset D not only provides a simpler way to represent the data and ease the interpretation of the results, but also retains all the relevant information for the classification providing some of the highest AUC values. For that reasons cS1-cS2 is the best data representation. The only drawback is the determination of the correction factors, which implies a huge experimental effort. In that sense, dataset C3 composed of only five features that can be extracted directly from the raw signals, can be useful to cross-check the modeling of those factors.
@@
Another important conclusion is that we have not found a significant difference in performance between the ML algorithms. Even more, the simple and very fast XGBoost method is usually on-par with more complex and harder to tune architectures. Additionally, it has the advantage of being more robust, in the sense that it is less sensitive to its hyperparameters, weight initialization, and yields a smaller uncertainty value. In Fig. 5 we show the ROC curves for some data representations and XGBoost, they correspond to ones with the lowest and highest AUC. The rest of the algorithms have similar ROC shapes, according to their AUC.